# LlamaTouch: A Faithful and Scalable Testbed for Mobile UI Automation Task Evaluation

### v1

@@ Abstract.
The emergent large language/multimodal models facilitate the evolution of mobile agents, especially in the task of mobile UI automation. However, existing evaluation approaches, which rely on human validation or established datasets to compare agent-predicted actions with predefined ones, are unscalable and unfaithful. To overcome these limitations, this paper presents LlamaTouch, a testbed for on-device agent execution and faithful, scalable agent evaluation. By observing that the task execution process only transfers UI states, LlamaTouch employs a novel evaluation approach that only assesses whether an agent traverses all manually annotated, essential application/system states. LlamaTouch comprises three key techniques: (1) On-device task execution that enables mobile agents to interact with real mobile environments for task completion. (2) Fine-grained UI component annotation that merges pixel-level screenshots and textual screen hierarchies to explicitly identify and precisely annotate essential UI components with a rich set of designed annotation primitives. (3) A multi-level state matching algorithm that utilizes exact and fuzzy matching to accurately detect critical information in each screen with unpredictable UI layout/content dynamics. LlamaTouch currently incorporates four mobile agents and 495 UI automation tasks, encompassing both tasks in the widely-used datasets and our self-constructed ones for more diverse mobile applications. Evaluation results demonstrate the LlamaTouch’s high faithfulness of evaluation in real environments and its better scalability than human validation. LlamaTouch also enables easy task annotation and integration of new mobile agents. Code and dataset are publicly available at https://github.com/LlamaTouch/LlamaTouch.
copyright: none
@@ 1. Introduction
Mobile intelligent agents empower users to interact with their smartphones using natural languages, alleviating them from tedious, cumbersome smartphone operations. These agents are particularly beneficial for individuals with visual or hand impairments, or in situations where using a screen is not practical (e.g., driving). Notable mobile agents, such as Apple Siri (Apple, 2024) and Google Assistant (Google, 2024b), have become indispensable services on smartphones. The recent advent of large language models (LLMs) and multimodal LLMs have facilitated researchers building more powerful mobile agents (Li et al., 2024; Wen et al., 2023; Wang et al., 2023; Lee et al., 2023). The key capability of those agents is to comprehend user instructions in natural language and execute corresponding actions on the mobile interface, as called mobile UI automation, e.g., “forward the last email from Bob to Alice”.
@@
Despite claims of powerful UI automation capabilities achieved by recent LLM-powered mobile agents, their evaluation methods are somewhat flawed. Unlike traditional machine learning models evaluated on well-established, static datasets, mobile agents need to interact with the dynamic, indeterministic states of a smartphone (e.g., network connectivity, dynamic content) as inputs. Simply evaluating mobile agents using deterministic smartphone states from datasets cannot uncover their true capabilities (Rawles et al., 2023; Li et al., 2020; Sun et al., 2022).
@@
In general, there are two methods to evaluate mobile UI automation tasks, but neither achieves both high faithfulness and scalability. (1) The most intuitive approach is to request humans to verify the completion of tasks. However, human evaluation is difficult to reproduce (Chiang and Lee, 2023). The requisite human effort increases with the number of agents, tasks, and evaluation platforms. (2) The most popular approach used in most prior arts (Rawles et al., 2023; Yan et al., 2023; Sun et al., 2022; Wen et al., 2023; Zhan and Zhang, 2023; Hong et al., 2023) is exact action match on established datasets, akin to traditional machine learning evaluations. The key idea is to ask annotators to generate a correct sequence of actions that succeed on the task as data labels, and then compare agent-generated actions to these labels. Although this approach allows for some error tolerance, e.g., the variations in click positions on the screen (Rawles et al., 2023), it cannot cover all possible and “infinite” paths to complete a UI automation task. Consequently, it leads to a significantly higher false negative rate. For example, for task “Reserve a rental car in Los Angeles from June 1st-7th, with a budget of up to \$60 per day on Expedia”, the sequence of three filtering actions can be interchanged. Only taking one of these execution paths as the reference may incorrectly verify a task that is essentially completed. Moreover, LLM-powered agents are known to be able to self-correct their wrong actions (Pan et al., 2023), which is critical to enhancing UI automation capabilities, yet is impossible to evaluate in a static dataset. These limitations are further demonstrated in $\S$§$2.$
@@
This paper presents LlamaTouch, the first testbed for evaluating mobile agents on real-world mobile environments without compromising faithfulness and scalability. The key idea of LlamaTouch is to check the task execution trace against a few “essential states” identified by the annotators, rather than matching them against predefined action sequences in static traces. For instance, the essential states for the task “open app Microsoft Excel (install if not already installed), go to login, …” should include (1) the application “Microsoft Excel” is opened; and (2) the application is on the login page. Other operations, like app installation, are considered non-essential and should be ignored. During task execution, LlamaTouch enables mobile agents to retrieve only the task description from the static dataset, while device states are directly acquired from real-world devices. Actions produced by mobile agents are executed on those smartphones, and all UI interaction data are recorded as task execution traces. In the evaluation phase, LlamaTouch compares task execution traces with annotated essential states to determine whether a task has been completed.
@@
To ensure faithful and scalable evaluation, LlamaTouch integrates two effective methods. (1) LlamaTouch adopts a fine-grained labeling mechanism for essential state annotation at both the screen level and single UI component level. It combines pixel-level screenshots and textual screen hierarchies to explicitly highlight important UI components. With a rich set of annotation primitives provided by LlamaTouch, it reduces human efforts to heuristically identify and annotate the attributes of essential states for evaluation, e.g., the content of a textbox should be exactly matched. These annotated primitives are subsequently used for faithful evaluation. (2) During the evaluation process, LlamaTouch employs a multi-level state match algorithm that combines fuzzy and exact matches on diverse annotated UI states. It uses (i) approximate screen match that enables LlamaTouch to adapt to dynamic mobile environments and varying screen contents, and (ii) mixed UI state match that detects and matches critical on-screen information.
@@
Dataset and testbed. We present a large-scale dataset with pre-annotated essential states for evaluating mobile UI automation tasks in real-world Android environments. This dataset includes 495 distinct tasks encompassing a wide array of popular Android applications. We complement this dataset with an easy-to-use testbed that enables mobile agents to interact seamlessly with real Android environments. This testbed provides a collection of concise, widely used APIs, ensuring compatibility with most mobile agents. Mobile agents can be easily integrated into LlamaTouch and use our dataset to test their capabilities in mobile UI automation tasks in real-world scenarios.
@@
Evaluation. We implemented LlamaTouch by utilizing Google Android emulator (Google, 2023) as actual smartphones. Currently, LlamaTouch has four built-in agents, AutoUI (Zhan and Zhang, 2023), AppAgent (Yang et al., 2023), AutoDroid (Wen et al., 2023), and CoCo-Agent (Ma et al., 2024), along with 495 diverse tasks. The number of supported agents and tasks will be extended in the future. With human validation as the ground truth for task completion, LlamaTouch achieves nearly 80% accuracy in detecting completed tasks in real-world environments, while exact action match on static datasets fails to do so. We also reveal the limitations of current mobile agents in handling tasks practically in real-world environments.
@@
  • We observed the weakness of high false negative rates in evaluating mobile UI automation agents using static datasets. To address this, we proposed an evaluation design that only compares essential states rather than concrete action sequences.
@@
  • We devised a method for annotating essential states using a variety of annotation primitives. This approach combines visually intuitive screenshots with semantically precise VHs to enable fine-grained and accurate UI component localization and annotation.
@@
  • We designed a novel task evaluation system that employs both exact and fuzzy matching at various UI state levels. It enables faithful evaluation of mobile agents and scales well to dynamic execution environments.
@@
  • We proposed LlamaTouch, the first testbed to faithfully and salably evaluate mobile UI automation agents in real-world mobile environments. It comprises 495 UI automation tasks with human-annotated essential states. Four agents integrated in LlamaTouch demonstrate its faithfulness and scalability in UI automation task evaluation.
2. Background and Motivation
2.1. Agents for Mobile UI Automation
Mobile agents have simplified the cumbersome and dull operations on smartphones for users. The progression of mobile agents for mobile UI task automation can be categorized into three phases. (1) API-based agents like Google Assistant (Google, 2024b) and Apple Siri (Apple, 2024) interact with applications through pre-defined application programming interfaces. This approach is reliable while limited in structured and predictable tasks. (2) Learning-based agents (Rawles et al., 2023; Sun et al., 2022; Li et al., 2020; Zhan and Zhang, 2023) utilize deep learning techniques to learn from previous mobile interaction traces, but their capabilities were still confined by their training data. (3) Recently, LLMs and multi-modal LLMs have revolutionized the capabilities of mobile agents (OpenAI, 2023; Hong et al., 2023; Wen et al., 2023; Wang et al., 2023). These models, owing to their vast knowledge base, can understand complex, real-world scenarios. Mobile agents powered by these models are able to interpret natural language instructions accurately and translate them into actionable tasks on smartphones’ screens. This evolution marks a significant leap in the flexibility and adaptability of mobile agents.
Mobile agents typically operate with the following components. Controller is the brain of mobile agents. It interprets task instructions and UI contexts, and generates actions to be executed on the current UI context. The widely-used controllers include DL models tailored for specific applications (Rawles et al., 2023; Sun et al., 2022; Li et al., 2020; Zhan and Zhang, 2023), LLMs (GPT-3.5, GPT-4) (Wen et al., 2023; Wang et al., 2023; Lee et al., 2023), and multi-modal LLMs (GPT-4V) (Yan et al., 2023; Hong et al., 2023; Yang et al., 2023).
Input: UI representation. Existing mobile agents take a task description and a UI representation as the input of their controller. There are two basic types of UI representations: screenshots and VH. A screenshot is a visual capture of the current screen. A text-format VH provides a tree-like structure of the UI elements present on a screen, including their properties such as type, position, and text contents. On top of screenshots and VH, some controllers further extract UI semantics to enhance UI understanding. For example, Yan et al. (Yan et al., 2023) overlies numeric tags on top of each text and icon detected by OCR tools; AXNav (Taeb et al., 2023) also converts screenshots to bounding boxes and labels, enabling them comprehensible by LLMs. Further processing based on VH, like converting it to simple HTML representations, is also widely utilized (Wang et al., 2023; Wen et al., 2023).
Output: actions. The output of controllers are actions to be executed on the current screen, such as click, swipe, and input text. Action parameters can be abstracted at different levels depending on the agent’s design and their input format. (1) Concrete coordinates on the screen (Rawles et al., 2023; Sun et al., 2022; Hong et al., 2023; Zhan and Zhang, 2023). It operates as a direct interaction with the screen like human operations. (2) Icon marker (Yan et al., 2023; Yang et al., 2023). The output target will specify a specific icon or graphical element within the UI representation. (3) HTML index (Wang et al., 2023; Wen et al., 2023; Lee et al., 2023). By ingesting HTML representations, controllers will give a concrete HTML index as the action target, which matches specific elements (icons or text) on the screen.
@@ 2.1. Agents for Mobile UI Automation
Table 1. The comparison between mobile agent benchmarks. LlamaTouch is the first testbed designed for mobile agents driven by essential state match. LlamaTouch also supports fine-grained UI-guided essential state annotation with a rich set of primitives covering a wide array of diverse UI states.
@@
Figure 1. Two major limitations of evaluating mobile agents on static datasets.
$2.$ Mobile UI Automation Benchmarks
@@
As shown in Table 1, a variety of datasets and environments are proposed to evaluate mobile agents in UI automation tasks, while none of them achieve both faithfulness and scalability.
@@
Some work such as Rico (Deka et al., 2017), PixelHelp (Li et al., 2020), and AITW (Rawles et al., 2023), provides static datasets (task descriptions and UI representations). Mobile agents predict concrete actions on static UI representations, which will be compared with the ground-truth actions in datasets. While this approach is straightforward, it is insufficient to reveal the performance of mobile agents for two reasons. (1) Inaccurate exact action match are pervasive when a functionally correct action is deemed incorrect due to different action parameters. Figure 1(a) shows two cases. First, for click actions, the clickable area defined in the dataset may be narrow. While in reality, the area might encompass the entire UI component (marked with the red bounding box). Second, non-identical text inputs can lead to the same correct result (marked with green bounding boxes) in most search tasks. This issue has also been observed in previous literature but has not been solved (Yan et al., 2023; Rawles et al., 2023). (2) Lack of tolerance for different execution paths. In real-world environments, an UI automation task can usually be completed in various paths, as demonstrated in Figure 1(b1) and Figure 1(b2). However, pre-defined datasets might only provide one deterministic path for reference, leading to inaccurate evaluation.
@@
There are also work that enable agent execution in real-world environments. AndroidEnv (Toyama et al., 2021) and Mobile-Env (Zhang et al., 2023) enable the execution of mobile automation tasks in real environments. However, they do not essentially support essential state match during the end-to-end task execution, therefore compromising evaluation accuracy. AndroidArena (Xing et al., 2024) observed the weakness of step-wise action match on static datasets does not fully tolerant redundant actions in task execution paths. They proposed a subsequence-based action match: a task is treated as completed when it contains the ground-truth action sequence as its subsequence. We take AndroidArena as a baseline in $\S$§4.3 to compare its evaluation accuracy with LlamaTouch. WebArena (Zhou et al., 2023) provides a real playground for Web agents. It uses essential states to evaluate task completion (e.g., the final result should be or should include some key information). LlamaTouch differs from WebArena on the mobile platform in both the essential state annotation and evaluation process: (i) LlamaTouch combines the visual screenshot with textual view hierarchy of the same screen for fine-grained and precise UI component identification; (ii) LlamaTouch uses a richer set of primitives to comprehensively annotate essential UI states and faithfully evaluate them even with high screen content dynamics.
@@
Human validation is usually used to validate whether a UI automation task is completed (Rawles et al., 2023; Taeb et al., 2023). However, the cost of human validation is too high, thereby scaling poorly to multiple tasks, agents, and mobile devices. LlamaTouch ensures high scalability as evaluating on static datasets while preserving faithfulness as human validation.
Figure 2. The workflow of on-device task execution.

### v2

# LlamaTouch: A Faithful and Scalable Testbed for Mobile UI Task Automation

@@ Abstract.
The emergent large language/multimodal models facilitate the evolution of mobile agents, especially in mobile UI task automation. However, existing evaluation approaches, which rely on human validation or established datasets to compare agent-predicted actions with predefined action sequences, are unscalable and unfaithful. To overcome these limitations, this paper presents LlamaTouch, a testbed for on-device mobile UI task execution and faithful, scalable task evaluation. By observing that the task execution process only transfers UI states, LlamaTouch employs a novel evaluation approach that only assesses whether an agent traverses all manually annotated, essential application/system states. LlamaTouch comprises three key techniques: (1) On-device task execution that enables mobile agents to interact with realistic mobile environments for task execution. (2) Fine-grained UI component annotation that merges pixel-level screenshots and textual screen hierarchies to explicitly identify and precisely annotate essential UI components with a rich set of designed annotation primitives. (3) A multi-level application state matching algorithm that utilizes exact and fuzzy matching to accurately detect critical information in each screen, even with unpredictable UI layout/content dynamics. LlamaTouch currently incorporates four mobile agents and 496 tasks, encompassing both tasks in the widely-used datasets and our self-constructed ones to cover more diverse mobile applications. Evaluation results demonstrate LlamaTouch’s high faithfulness of evaluation in real-world mobile environments and its better scalability than human validation. LlamaTouch also enables easy task annotation and integration of new mobile agents. Code and dataset are publicly available at https://github.com/LlamaTouch/LlamaTouch.
mobile agent, UI task automation, evaluation, testbed
journalyear: 2024copyright: acmlicensedconference: The 37th Annual ACM Symposium on User Interface Software and Technology; October 13–16, 2024; Pittsburgh, PA, USAbooktitle: The 37th Annual ACM Symposium on User Interface Software and Technology (UIST ’24), October 13–16, 2024, Pittsburgh, PA, USAdoi: 10.1145/3654777.3676382isbn: 979-8-4007-0628-8/24/10ccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
Mobile intelligent agents empower users to interact with their smartphones using natural languages, alleviating them from tedious and cumbersome smartphone operations. These agents are particularly beneficial for individuals with visual or hand impairments, or in situations where using a screen is not practical (e.g., driving). Notable mobile agents, such as Apple Siri (Apple, 2024) and Google Assistant (Google, 2024b), have become indispensable services on smartphones. The recent advent of large language models (LLMs) and multimodal LLMs has facilitated researchers in building more powerful mobile agents (Li et al., 2024; Wen et al., 2024; Wang et al., 2023; Lee et al., 2023). The key capability of these agents is to comprehend user instructions in natural language and execute corresponding actions on the mobile interface, as called mobile UI task automation, e.g., “forward the last email from Bob to Alice”.
@@
Despite claims of powerful task automation capabilities achieved by recent LLM-powered mobile agents, their evaluation methods are somewhat flawed. Unlike traditional machine learning models evaluated on well-established static datasets, mobile agents need to interact with the dynamic and indeterministic states of a smartphone (e.g., network connectivity, dynamic content) as inputs. Additionally, mobile devices use touch and gesture-based interactions (e.g., swipe, pinch), leading to diverse and ambiguous inputs. This variability complicates static action-based matching algorithms. Therefore, simply evaluating mobile agents using deterministic smartphone states from datasets cannot uncover their true capabilities (Rawles et al., 2023; Li et al., 2020; Sun et al., 2022).
@@
In general, there are two methods to evaluate mobile UI automation tasks, but neither achieves both high faithfulness and scalability. (1) The most intuitive approach is to request humans to verify the completion of tasks. However, human evaluation is difficult to reproduce (Chiang and Lee, 2023), and the requisite human effort increases with the number of agents, tasks, and evaluation platforms. (2) The most popular approach used in most prior work (Rawles et al., 2023; Yan et al., 2023; Sun et al., 2022; Wen et al., 2024; Zhan and Zhang, 2023; Hong et al., 2023) is exact action match on established datasets, akin to traditional machine learning evaluations. The key idea is to ask annotators to generate a correct sequence of actions that succeed on the task as data labels, and then compare agent-generated actions to these labels. Although this approach allows for some error tolerance, e.g., the variations in click positions on the screen (Rawles et al., 2023), it cannot cover all possible and “infinite” paths to complete a UI automation task. Consequently, it leads to a significantly higher false negative rate. For example, for the task “Reserve a rental car in Los Angeles from June 1st-7th, with a budget of up to \$60 per day on Expedia”, the sequence of three filtering actions can be interchanged. Only taking one of these execution paths as the reference may incorrectly verify a task that is essentially completed. Moreover, LLM-powered agents are known to be able to self-correct their wrong actions (Pan et al., 2023), which is critical to enhancing UI task automation capabilities, yet is impossible to evaluate in a static dataset. These limitations are further demonstrated in §$\S$2.2.
@@
This paper presents LlamaTouch, the first testbed for evaluating mobile agents in real-world mobile environments without compromising faithfulness and scalability. The key idea of LlamaTouch is to check the task execution trace against a few “essential states” identified by the annotators, rather than matching them against predefined action sequences in static traces. For instance, the essential states for the task “open app Microsoft Excel (install if not already installed), go to login, …” should include (1) the application “Microsoft Excel” is opened, and (2) the application is on the login page. Other operations, like app installation, are considered non-essential and should be ignored. During task execution, LlamaTouch enables mobile agents to retrieve only task descriptions from static datasets, while device states are directly acquired from realistic mobile devices. Actions produced by mobile agents are directly operated on those devices, and all UI interaction data are recorded as task execution traces. In the evaluation phase, LlamaTouch compares task execution traces with annotated essential states to determine whether a task has been completed.
@@
To ensure faithful and scalable evaluation, LlamaTouch integrates two effective methods. (1) LlamaTouch adopts a fine-grained labeling mechanism for essential state annotation at both the screen level and single UI component level. It combines pixel-level screenshots and textual screen hierarchies to explicitly highlight important UI components. With a rich set of annotation primitives provided by LlamaTouch, it reduces human effort to heuristically identify and annotate the attributes of essential states for evaluation, e.g., the text inside a textbox should be exactly matched. These annotated UI states are subsequently used for faithful evaluation. (2) During evaluation, LlamaTouch employs a multi-level state matching algorithm that combines fuzzy and exact matches on diverse annotated UI states. It uses (i) approximate screen matching, which enables LlamaTouch to adapt to dynamic mobile environments and varying screen contents, and (ii) mixed UI state matching, which detects and matches critical on-screen information.
@@
Dataset and testbed. We present a large-scale dataset with pre-annotated essential states for evaluating mobile UI automation tasks in real-world mobile environments. This dataset includes 496 distinct tasks encompassing a wide array of popular Android applications. We complement this dataset with an easy-to-use testbed that enables mobile agents to interact seamlessly with realistic Android environments. This testbed provides a collection of concise, widely used APIs, ensuring compatibility with most mobile agents. Mobile agents can be easily integrated into LlamaTouch and use our dataset to test their capabilities in mobile UI task automation in real-world scenarios.
@@
Evaluation. We implemented LlamaTouch by utilizing Google Android emulator (Google, 2023) and one Google Pixel 5 smartphone as realistic Android environments. Currently, LlamaTouch has four built-in agents, including AutoUI (Zhan and Zhang, 2023), AppAgent (Yang et al., 2023), AutoDroid (Wen et al., 2024), and CoCo-Agent (Ma et al., 2024), along with 496 diverse tasks. With human validation results as the ground truth for task completion, LlamaTouch achieves nearly 80% evaluation accuracy in detecting completed tasks in real-world environments, while prior action-based evaluation methods fail to do so. We also reveal the limitations of current mobile agents in handling tasks practically in real-world environments.
@@
  • We observed the weakness of high false negative rates in evaluating mobile UI task automation agents using static datasets. To address this, we proposed an evaluation design that only compares essential states rather than concrete action sequences.
@@
  • We devised a method for annotating essential states using a variety of annotation primitives. This approach combines visually intuitive screenshots with semantically precise view hierarchies to enable fine-grained and accurate UI component localization and annotation.
@@
  • We designed a novel task evaluation approach that employs both exact and fuzzy matching at various UI state levels. It enables faithful evaluation of mobile agents and adapts well to dynamic execution environments.
@@
  • We proposed LlamaTouch, the first testbed to faithfully and salably evaluate mobile UI task automation agents in real-world mobile environments. It comprises 496 tasks with human-annotated essential states. Four agents integrated in LlamaTouch demonstrate its faithfulness and scalability in UI automation task evaluation.
@@
Table 1. The comparison between mobile agent benchmarks. LlamaTouch is the first testbed designed for mobile agents driven by essential state matching. LlamaTouch also supports fine-grained UI-guided essential state annotation with a rich set of primitives covering a wide array of matching implementations.
@@
2. Background and Motivation
2.1. Agents for Mobile UI Task Automation
Mobile agents have simplified the cumbersome and dull operations on smartphones for users. The progression of mobile agents for mobile UI task automation can be categorized into three phases. (1) API-based agents like Google Assistant (Google, 2024b) and Apple Siri (Apple, 2024) interact with applications through predefined application programming interfaces. This approach is reliable while limited in structured and predictable tasks. (2) Learning-based agents (Rawles et al., 2023; Sun et al., 2022; Li et al., 2020; Zhan and Zhang, 2023) utilize deep learning techniques to learn from previous mobile interaction traces, but their capabilities are still confined by their training data. (3) Recently, LLMs and multi-modality LLMs have revolutionized the capabilities of mobile agents (OpenAI, 2023; Hong et al., 2023; Wen et al., 2024; Wang et al., 2023). These models, owing to their vast knowledge base, can understand complex, real-world mobile screens. Mobile agents powered by these models can accurately interpret natural language instructions and translate them into actionable tasks on smartphone screens. This evolution marks a significant leap in the flexibility and adaptability of mobile agents.
Mobile UI task automation agents typically operate with the following components.
Controller is the brain of mobile agents. It interprets task instructions and UI contexts, and then generates actions to be executed on the current UI context. Widely-used controllers include deep learning models tailored for specific applications (Rawles et al., 2023; Sun et al., 2022; Li et al., 2020; Zhan and Zhang, 2023), LLMs (e.g., GPT-4, Llama) (Wen et al., 2024; Wang et al., 2023; Lee et al., 2023), and multi-modality LLMs (e.g., GPT-4V) (Yan et al., 2023; Hong et al., 2023; Yang et al., 2023).
Input: UI Representation. Existing mobile agents take a task description and UI representations as the input of their controller. There are two basic types of UI representations: screenshot and view hierarchy (VH). A screenshot is a visual capture of the current screen. A VH provides a textual tree-like structure of the UI elements present on a screen, including their properties such as type, position, and text contents. On top of screenshots and VHs, some controllers further extract UI semantics to enhance UI understanding. For example, Yan et al. (Yan et al., 2023) overlay numeric tags on top of each text and icon detected by OCR tools; AXNav (Taeb et al., 2023) converts screenshots to bounding boxes and labels, making them comprehensible to LLMs. Further processing based on VH, such as converting it to simple HTML representations, is also widely utilized (Wang et al., 2023; Wen et al., 2024).
Output: Action. The output of controllers consists of actions to be executed on the current screen, such as click, swipe, and input text. Action parameters can be abstracted at different levels depending on the agent’s design and input format. (1) Concrete coordinates on the screen (Rawles et al., 2023; Sun et al., 2022; Hong et al., 2023; Zhan and Zhang, 2023): This operates as a direct interaction with the screen, similar to human operations. (2) Icon marker (Yan et al., 2023; Yang et al., 2023): The output target will specify a specific icon or graphical element within the UI representation. (3) HTML index (Wang et al., 2023; Wen et al., 2024; Lee et al., 2023): By ingesting HTML representations, controllers will give a concrete HTML index as the action target, which matches specific elements (icons or text) on the screen.
(a) Inaccurate action match in two tasks. Failed reasons: “Left”: wrong click parameter; “Right”: wrong input text.
(b) Different task execution paths lead to the same screen.
@@ 2.1. Agents for Mobile UI Task Automation
Figure 1. Two major limitations of evaluating mobile UI task automation on static datasets.
2.2. Mobile UI Task Automation Benchmarks
Figure 2. LlamaTouch workflow and code demonstrations for mobile UI task execution and trace evaluation. LlamaTouch enables mobile UI task automation agents to integrate easily with AgentEnv for on-device task execution with minimal programming effort. Agent execution traces recorded by AgentEnv are used in conjunction with the LlamaTouch dataset in a separate evaluation process.
@@ 2.2. Mobile UI Task Automation Benchmarks
As shown in Table 1, a variety of datasets and environments are proposed to evaluate mobile agents in UI task automation, but none of them achieve both faithfulness and scalability.
@@
Some work such as Rico (Deka et al., 2017), PixelHelp (Li et al., 2020), and AITW (Rawles et al., 2023), provides static datasets with task descriptions, UI representations, and actions sequences. Mobile agents predict concrete actions on static UI representations, which are compared with the ground-truth actions in the datasets. While this approach is straightforward, it is insufficient to reveal the performance of mobile agents for two reasons. (1) Inaccurate exact action match. Functionally correct actions may be deemed incorrect due to different action parameters. Figure 1 illustrates two cases. First, for click actions, the clickable area defined in the dataset may be narrow, whereas in reality, the area might encompass the entire UI component (marked with the red bounding box). Second, non-identical text inputs can lead to the same correct search result (marked with green bounding boxes) in most search tasks. This issue has been observed in previous literature but remains unsolved (Yan et al., 2023; Rawles et al., 2023). (2) Lack of tolerance for different execution paths. In real-world environments, a task can usually be completed in various paths based on different device/application states, as shown in Figure 1. However, predefined datasets might only provide one deterministic path for reference, leading to inaccurate evaluation.
@@
There is also other work that enables agent execution in real-world environments, such as AndroidEnv (Toyama et al., 2021) and Mobile-Env (Zhang et al., 2023). However, they do not inherently support essential state match during end-to-end task execution, therefore compromising evaluation accuracy. AndroidArena (Xing et al., 2024) observed the weakness of step-wise action match on static datasets: it does not fully tolerate redundant actions in task execution paths. They proposed a subsequence-based action match, where a task is treated as completed if it contains the ground-truth action sequence as its subsequence. We take AndroidArena as a baseline in §$\S$5.3 to compare its evaluation accuracy with LlamaTouch. WebArena (Zhou et al., 2023) provides a realistic playground for web agents. It uses essential states to evaluate task completion (e.g., the final result should be or should include some key information). LlamaTouch differs from WebArena on the mobile platform in both essential state annotation and evaluation process: (i) LlamaTouch combines visual screenshots with textual VHs of the same screens for fine-grained and precise UI component identification. (ii) LlamaTouch uses a richer set of primitives to comprehensively annotate essential UI states and faithfully evaluate them even with high screen content dynamics.
@@
Human validation is usually used to validate whether a UI automation task is completed (Rawles et al., 2023; Taeb et al., 2023). However, the cost of human validation is too high, scaling poorly to multiple tasks, agents, and mobile devices. LlamaTouch ensures high scalability as with evaluating on static datasets while preserving faithfulness similar to human validation.